OpenStudy (anonymous):

Find an equation of the line parallel to 6x+2y=49 containing the point (4,9). Put your equation in slope-intercept form.

OpenStudy (anonymous):

Parallel lines have the same slope.

OpenStudy (anonymous):

Could you give me the equation using the point (4,9)?

OpenStudy (anonymous):

First you need to determine the slope of the parallel line 6x+2y=49

OpenStudy (anonymous):

Could you show me?

OpenStudy (anonymous):

Slope formula y=mx+c. Slope = m.

OpenStudy (anonymous):

Could your work this problem out so I can take it to a similar one and see if I got it? Thanks. I learn a lot more from seeing one done and doing another.

OpenStudy (anonymous):

EXample 4x+2y = 24..thus 2y=-4x+24 as I subtracted 4x from both sides of the equation..now we have 2y=-4x+24.....Now I divide both sides of the equation by 2 and we end up with 2(Y)/2=-4/2(X) +24/2 or y=-2x+12....thus the slope of this line is -2.

OpenStudy (anonymous):

So what would you do then?